# PRISM: Proof-Carrying Artifact Generation through LLM $\times$ MDE Synergy and Stratified Constraints

### v1

@@ 1. Stage 1:
    Generation with real-time constraint enforcement. Layer-1 constraints are compiled into a prefix-closed automaton (deterministic finite automaton for regular structures or bounded pushdown automaton for context-free patterns, unfolded to depth $d$; see Section 3.4). During decoding, at each step $t$, the executor computes the allowed token set $M_{t}=\{y\in\Sigma:\delta(s_{t},y)\neq\bot\}$, where $s_{t}$ is the current automaton state and $\delta$ is the transition function. Invalid continuations are masked before the LLM samples the next token. This guarantees prefix safety: every partial output always remains on a path that could complete to a structurally valid artifact (Theorem 12). In parallel, PRISM records an audit trail $\tau_{t}=\langle s_{t},M_{t},y_{t},\delta(s_{t},y_{t}),\Delta t\rangle$, which captures the control decisions that led to each generated token for later auditing.
@@ 2. Stage 2:
    Post-generation formal verification. After the candidate artifact is produced, PRISM runs Layer-2 validators. Semantic validators check constraints such as cross-file reference integrity and role consistency; logic validators check temporal and numeric properties such as ordering constraints and resource bounds. Each validator emits a machine-checkable certificate $\pi_{\bullet}\in\{\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}}\}$ (e.g., DFA acceptance traces, semantic validation reports, logic proofs/unsat cores). Crucially, these certificates are independent of the generator: an external auditor can re-check correctness using only the artifact $a$, the evidence bundle $\Pi$, and the validator identifiers—without having to trust the LLM itself (Definition 13).
@@ 3. Stage 3:
    Audit-Guided Repair (AGR). If validation fails, PRISM does not blindly re-generate from scratch. Instead, AGR analyzes the composite evidence $\Pi=(\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}})\oplus\tau$ to identify the root cause. Structural violations are localized using the DFA state in $\tau$ (e.g., which mandatory element was skipped); semantic violations are localized using violation paths from semantic validators (e.g., which reference is dangling); logical violations are localized using unsat cores from logic solvers (e.g., which constraints are mutually inconsistent) (Algorithm 4). AGR then invokes the LLM in a constrained repair mode: the prompt includes the violation diagnostics plus the relevant UMM/ICM context, guiding the model to propose targeted fixes rather than producing an unrelated new artifact.
@@
Figure 1 summarizes how these components interact. Natural-language requirements are first interpreted by an LLM-based NLP Processing Layer, while explicit domain schemas are parsed by a Knowledge-Graph Construction pipeline that builds the UMM and populates the ICM through Channel 1 and Channel 2. The Unified Constraint Layer then compiles these constraints into L1 automata and L2 validators. During Instance Model Generation, the LLM receives retrieval-augmented UMM subgraphs as context, and the Structure Decoder enforces L1 constraints through token masking while logging audit tuples $\tau_{t}$ to an Evidence Registry. After generation, the Validation & Repair Layer executes L2 validators (structural $\rightarrow$ semantic $\rightarrow$ logical), appends each certificate $\pi_{\bullet}$ to the registry, and either: (i) assembles the final verifiable artifact $\mathfrak{A}=(a,\Pi,\varphi)$ by hashing the artifact, the evidence, and a timestamp $H(a\|\Pi\|t)$; or (ii) triggers AGR to request a focused repair from the LLM or, in edge cases, a manual override by a domain expert. This “graduated automation” design lets PRISM automate routine generation while keeping human reviewers in control of exceptional cases, now equipped with machine-checkable diagnostics instead of ad hoc debugging.
Roadmap for the Remainder of the Section.
The following subsections elaborate each component. Section 3.2 details the construction of the UMM and explains how heterogeneous domain specifications are merged into a single auditable model. Section 3.3 describes how the ICM unifies deductive constraints (Channel 1) and LLM-assisted constraints (Channel 2) under the UMM, including entity linking and consistency checks. Section 3.4 presents CVG, including automaton-guided decoding (L1) and semantic/logic validation (L2), and proves prefix safety under bounded unfolding. Section 3.5 defines the composite evidence $\Pi$ and explains how PRISM packages artifacts together with machine-checkable certificates. Section 3.6 introduces the audit trail recorder, the evidence registry, and the AGR loop, and analyzes repair convergence in terms of layered violation dynamics.
@@ 3.2. Unified Meta-Model Construction: A Dual-Path Approach
In practice, safety-critical domains present two very different starting points for model construction. Some domains (e.g., AUTOSAR, AADL, OPC UA) already publish an explicit meta-model: classes, attributes, and relationships are described in machine-readable form (XSD/XMI/Ecore) or in tightly maintained technical documentation. Other domains (e.g., regulatory and legal workflows, device compliance manuals) do not provide such an explicit schema, even though they implicitly assume a fairly well-structured universe of entities (“who is responsible,” “what object is configured,” “what reference must exist”). PRISM introduces the Unified Meta-Model (UMM) as the common semantic backbone across both cases. When an explicit meta-model exists, UMM ingests and normalizes it; when it does not, UMM inductively assembles the same kind of typed entity–relation structure from natural-language specifications with LLM assistance. In both cases, the result is a single, provenance-aware representation of the domain’s core entities and relationships that downstream components can treat as the source of truth. Its construction must satisfy three goals: (i) reconcile heterogeneous sources into one typed graph, (ii) maintain provenance so each element in the UMM can be traced back to its origin, and (iii) expose a stable structure that downstream components can rely on for generation, validation, and repair.
@@
Given a collection of such graphs $\{M_{i}=(V_{i},E_{i},T_{i},\mathcal{C}_{i})\}_{i\in I}$, the Unified Meta-Model $M_{\text{UMM}}$ is obtained by merging them into a single typed knowledge structure. In our implementation, this merge follows a colimit-style typed graph unification in the sense of (Diskin and Maibaum, 2012): overlapping entities and relations across different $M_{i}$ are identified and merged through structure-preserving homomorphisms. This provides two desirable properties for PRISM: (1) semantic coherence—concepts that represent the same domain notion across sources are aligned rather than duplicated; and (2) auditability—every node and edge in $M_{\text{UMM}}$ retains provenance links to its source model(s), allowing downstream validators and human reviewers to trace “where this concept came from.”
Figure 2. Two ways to build the Unified Meta-Model (UMM). Path S1 handles structured sources: when a machine-readable meta-model (for example AUTOSAR, AADL, OPC UA) already exists, PRISM ingests and normalizes it. Path S2 handles unstructured sources: when only natural-language requirements are available, PRISM uses an LLM to induce domain entities and relations directly from text. Deployments typically follow one of these two paths, depending on which input is available.:
@@
PRISM supports two alternative construction paths for the UMM (Figure 2). In practice, deployments usually select one of these paths: either the domain already ships a machine-readable meta-model (Path S1), or it only has natural-language specifications with no canonical XSD/UML (Path S2). Both paths produce the same kind of output: a typed graph $M_{\text{UMM}}$ with provenance annotations. We describe the two paths below.
@@
When an explicit domain meta-model exists, PRISM reuses an established dual-stage transformation process (Ma et al., 2025a). This process proceeds as follows:
@@
Path S1 therefore produces a UMM instance in which: (i) entities, attributes, and relations reflect the authoritative schema; and (ii) selected behavioral expectations are explicitly attached to those entities. Within PRISM, this serves as the canonical semantic backbone whenever the domain already provides a formalized meta-model.
@@
Stage 1: Context-Enriched Document Segmentation. A corpus $D=\{d_{1},\ldots,d_{n}\}$ of natural-language specifications is segmented by a learned boundary detector $B$, which identifies sections likely to define entities, attributes, or relationships. For each segment $s_{i}$, PRISM injects domain context into the prompt (e.g., known entity names, typical field names, role labels from prior segments or from a seed ontology). This grounding step constrains the extractor’s vocabulary and reduces hallucinated concepts.
@@
where $\alpha(c)=\bot$ indicates that $c$ does not match any known node. If $\alpha(c)\neq\bot$, the candidate is treated as an alias or refinement of the matched node, and its provenance metadata is attached to that node. If $\alpha(c)=\bot$, PRISM tentatively adds a new node or relation into $M_{\text{UMM}}$, tagged with provenance and marked for optional human review when the stakes are high (e.g., safety-critical components). Crucially, at this stage we are only inducing and reconciling domain entities, attributes, and relations; constraint extraction and formal consistency checking are handled later by the Integrated Constraint Model (ICM) in Section 3.3. No SHACL or SMT reasoning is applied here.
@@
Although UMM construction is primarily an integration activity rather than a new theoretical contribution, its runtime characteristics matter for practical deployment. Let $n_{i}=|V_{i}|$ be the number of entities extracted from source $M_{i}$, and let $m$ be the number of detected overlaps across all sources. The reconciliation procedure maintains disjoint sets of candidate-equivalent nodes and performs unions as they are aligned. Its complexity is summarized below.
Lemma 2 (Construction Complexity).
The dual-path construction algorithm (including node alignment and union over overlaps) runs in
(3) $\mathcal{O}\!\biggl(\textstyle\sum_{i}n_{i}+m\,\alpha\!\bigl(\sum_{i}n_{i}\bigr)\biggr)$
time and
(4) $\mathcal{O}\!\bigl(\textstyle\sum_{i}n_{i}\bigr)$
space, where $\alpha$ is the inverse Ackermann function from disjoint-set analysis.
Sketch.
Candidate overlaps are proposed via bounded subgraph matching to identify entities that are likely semantically identical. These candidates are merged using union-find with path compression. Union-find yields near-constant amortized merge cost, which gives the stated near-linear bounds in practice. ∎
The near-linear scaling implied by Lemma 2 allows PRISM to ingest industrial meta-models with thousands of entities without incurring the $O(n^{2})$ blow-up typical of manual pairwise weaving.
@@
Classical model weaving frameworks (e.g., AMW) rely on manually authored correspondence models, while graph transformation systems (e.g., AGG) require hand-written rules for every pair of source models. Purely LLM-driven extraction has also been explored (Chen et al., 2023), but without grounding the induced concepts in a persistent, provenance-tracked ontology, such approaches risk hallucinated entities and silently inconsistent references. PRISM differs in three respects:
  • Source Adaptability. Path S1 ingests explicit XSD/XMI-style meta-models through deterministic model transformations (Ma et al., 2025a). Path S2 induces a meta-model from prose using an LLM guided by domain context. Deployers choose whichever path matches their domain reality.
@@
  • Scalability. Because reconciliation proceeds via incremental union rather than exhaustive pairwise weaving, PRISM scales to realistic industrial domains, as summarized in Lemma 2.
In summary, UMM construction is the step that turns messy, heterogeneous domain knowledge into a single, provenance-aware semantic backbone. All later phases of PRISM—constraint modeling, generation-time enforcement, post-generation validation, and audit-guided repair—assume this backbone as the shared point of truth.
@@ 3.3. Constraint Extraction and Integrated Constraint Model (ICM)
The Unified Meta-Model (UMM) provides PRISM with a single, auditable semantic backbone. The Integrated Constraint Model (ICM) builds on that backbone by collecting, normalizing, and compiling the domain rules that generated artifacts must satisfy. These rules range from low-level structural requirements (e.g., mandatory fields, cardinalities) to high-level semantic and logical invariants (e.g., “an OperationInvokedEvent must reference an existing operation”), and they originate from both machine-readable meta-models and natural-language specifications. This subsection formalizes how such constraints are represented, how they are extracted through two complementary channels, how they are grounded in the UMM, and how they are compiled into executable Layer-1 (L1) and Layer-2 (L2) validators for generation-time enforcement and post-generation verification.
@@
Different inputs to PRISM (e.g., AUTOSAR XSD/XMI schemas vs. regulatory PDFs) induce different constraint spaces. Before integration, these must be semantically aligned with the UMM so that downstream validators can interpret them consistently.
@@
L2 Validator Instantiation.
The abstract semantic and logic validators $\mathcal{V}_{\mathrm{sem}}$ and $\mathcal{V}_{\mathrm{log}}$ in $\mathcal{I}$ can be instantiated through various verification formalisms depending on domain requirements. In our implementation, we instantiate semantic validators using SHACL (Shapes Constraint Language) (Knublauch and Kontokostas, 2017) for graph-structural constraints such as reference integrity, cardinality restrictions, and type compatibility. SHACL naturally captures cross-reference semantics in XML/RDF-based artifacts like AUTOSAR ARXML configurations. For logic validators, we employ SMT (Satisfiability Modulo Theories) solvers (de Moura and Bjørner, 2008), which efficiently handle bounded numeric reasoning, temporal constraints, and resource allocation predicates common in automotive timing specifications and legal precedence rules.
This instantiation choice reflects the structural characteristics of our evaluation domains (AUTOSAR, Brussels I bis). The validator interface remains extensible: alternative formalisms such as Alloy for relational constraints, OWL reasoners for ontology-based validation, or domain-specific checkers can substitute or complement SHACL/SMT where appropriate. The key architectural requirement is that validators produce machine-checkable certificates $\pi_{\text{sem}}$ and $\pi_{\text{logic}}$ with violation localization metadata sufficient to drive Audit-Guided Repair (Section 3.6).
@@
Channel 2: Inductive Extraction from Natural-Language Specifications. Many industrial domains encode mission-critical rules in prose rather than structured schemas: AUTOSAR PDFs, regulatory guidance, safety standards, legal jurisdiction rules. Channel 2 uses an LLM-based extractor to convert free-form text into structured constraint candidates, following prior work on LLMs as structured information extractors (Dagdelen et al., 2024; Lu et al., 2022; Wei et al., 2020; Jiao et al., 2024). Formally,
@@
Figure 3. LLM-based extraction from natural-language specifications. The extractor proposes structured constraint candidates and aligns each candidate to entities in the UMM. Only candidates that align cleanly and pass semantic checks are committed to the Integrated Constraint Model (ICM); the rest are flagged for human review instead of being turned directly into validators.:
@@
Figure 4. ICM as a multi-sorted algebra layered over the UMM. White nodes denote UMM entities; gray nodes encode semantic and logical constraints extracted from Channel 1 (deductive) and Channel 2 (LLM-assisted). Morphisms relate abstract constraints to class-/attribute-level constraints, enabling cross-layer reasoning and repair prioritization.:
@@
Algorithm 1 summarizes how PRISM assembles the ICM from both structured schemas and unstructured documents. This process is persistent: it populates the long-lived constraint repository (the ICM $\mathcal{I}$), rather than just synthesizing a one-off validator.
@@
In addition to the static ICM, PRISM supports request-specific constraints that apply only to a single generation run. For example, a user generating one AUTOSAR subsystem may require “all port connections in this subsystem must use compatible data types” without permanently altering the global repository. To accommodate such ad hoc requirements, we synthesize ephemeral L2 validators at runtime:
(13) $\mathrm{Synthesize}^{\mathrm{NLP}}:\mathcal{R}_{\text{dynamic}}\to\{\text{SHACL, SMT}\}.$
These validators are fed directly into L2 verification during CVG (Section 3.4), but they are not persisted in the ICM unless they pass solver-backed validation and are deemed reusable.

### v2

@@ 1. Stage 1:
    Generation with real-time constraint enforcement. Layer-1 constraints are compiled into a prefix-closed automaton (deterministic finite automaton for regular structures or bounded pushdown automaton for context-free patterns, unfolded to depth $d$; see Section 3.4). During decoding, at each step $t$, the executor computes the allowed token set $M_{t}=\{y\in\Sigma:\delta(s_{t},y)\neq\bot\}$, where $s_{t}$ is the current automaton state and $\delta$ is the transition function. Invalid continuations are masked before the LLM samples the next token. This guarantees prefix safety: every partial output always remains on a path that could complete to a structurally valid artifact (Theorem 10). In parallel, ATLAS records an audit trail $\tau_{t}=\langle s_{t},M_{t},y_{t},\delta(s_{t},y_{t}),\Delta t\rangle$, which captures the control decisions that led to each generated token for later auditing.
@@ 2. Stage 2:
    Post-generation formal verification. After the candidate artifact is produced, ATLAS runs Layer-2 validators. Semantic validators check constraints such as cross-file reference integrity and role consistency; logic validators check temporal and numeric properties such as ordering constraints and resource bounds. Each validator emits a machine-checkable certificate $\pi_{\bullet}\in\{\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}}\}$ (e.g., DFA acceptance traces, semantic validation reports, logic proofs/unsat cores). Crucially, these certificates are independent of the generator: an external auditor can re-check correctness using only the artifact $a$, the evidence bundle $\Pi$, and the validator identifiers—without having to trust the LLM itself (Definition 11).
@@ 3. Stage 3:
    Audit-Guided Repair (AGR). If validation fails, ATLAS does not blindly re-generate from scratch. Instead, AGR employs a validation-guided regeneration strategy. The framework analyzes the composite evidence $\Pi$ to localize violations: structural violations via DFA states in $\tau$, semantic violations via SHACL violation paths, and logical violations via SMT unsat cores (Algorithm 4). Unlike blind retries, AGR operates at file granularity but with constraint-aware guidance. It parses the diagnostics to identify specifically which constraints were violated and constructs a targeted repair prompt that explicitly injects these constraints as hard requirements. This narrows the search space and forces the LLM to attend to the specific failure modes (e.g., “missing reference” or “cardinality error”) that were missed in the first pass.
    The localized violations are then repaired through one of two routes:
    Route A: Automated validation-guided regeneration. For standard violations, AGR builds a constraint-aware repair prompt that explicitly injects the violated rules. The LLM performs a diagnostic-driven regeneration of the affected artifact, resolving the specific SHACL/SMT violation.
    Route B: Human-in-the-loop MDE editing. For complex semantic ambiguities, ATLAS projects the artifact into a context-aware configuration editor (automatically derived from the UMM, leveraging the structured transformation Path S1 detailed in Section 3.2). This transforms the human role from manual authoring to evidence-based decision making, allowing experts to resolve undecidable conflicts using machine-generated diagnostics.
@@
Figure 1 summarizes how these components interact. Natural-language requirements are first interpreted by an LLM-based NLP Processing Layer, while explicit domain schemas are parsed by a Knowledge-Graph Construction pipeline that builds the UMM and populates the ICM through Channel 1 and Channel 2. The Unified Constraint Layer then compiles these constraints into L1 automata and L2 validators. During Instance Model Generation, the LLM receives retrieval-augmented UMM subgraphs as context, and the Structure Decoder enforces L1 constraints through token masking while logging audit tuples $\tau_{t}$ to an Evidence Registry. The objective is to synthesize an instance model $a$ at the M1 layer that strictly conforms to the M2-level definitions in the UMM. After generation, the Validation & Repair Layer executes L2 validators (structural $\rightarrow$ semantic $\rightarrow$ logical), appends each certificate $\pi_{\bullet}$ to the registry, and either: (i) assembles the final verifiable artifact $\mathfrak{A}=(a,\Pi,\varphi)$ by hashing the artifact, the evidence, and a timestamp $H(a\|\Pi\|t)$; or (ii) triggers AGR to request a focused repair from the LLM or, in edge cases, a manual override by a domain expert. This “graduated automation” design lets ATLAS automate routine generation while keeping human reviewers in control of exceptional cases, now equipped with machine-checkable diagnostics instead of ad hoc debugging.
@@ 3.2. Unified Meta-Model Construction: A Dual-Path Approach
In practice, safety-critical domains present two very different starting points for model construction. Some domains (e.g., AUTOSAR, AADL, OPC UA) already publish an explicit meta-model: classes, attributes, and relationships are described in machine-readable form (XSD/XMI/Ecore) or in tightly maintained technical documentation. Other domains (e.g., regulatory and legal workflows, device compliance manuals) do not provide such an explicit schema, even though they implicitly assume a fairly well-structured universe of entities (“who is responsible,” “what object is configured,” “what reference must exist”). ATLAS introduces the Unified Meta-Model (UMM) as the common semantic backbone across both cases. When an explicit meta-model exists, UMM ingests and normalizes it; when it does not, UMM inductively assembles the same kind of typed entity–relation structure from natural-language specifications with LLM assistance. In both cases, the result is a single, provenance-aware representation of the domain’s core entities and relationships that downstream components can treat as the source of truth. Its construction must satisfy three goals: (i) reconcile heterogeneous sources into one typed graph, (ii) maintain provenance so each element in the UMM can be traced back to its origin, and (iii) expose a stable structure that downstream components can rely on for generation, validation, and repair.
@@
Given a collection of such graphs $\{M_{i}=(V_{i},E_{i},T_{i},\mathcal{C}_{i})\}_{i\in I}$, the Unified Meta-Model $M_{\text{UMM}}$ is obtained by merging them into a single typed knowledge structure. In our implementation, this merge follows a colimit-style typed graph unification in the sense of (diskin2012category): overlapping entities and relations across different $M_{i}$ are identified and merged through structure-preserving homomorphisms. This provides two desirable properties for ATLAS: (1) semantic coherence—concepts that represent the same domain notion across sources are aligned rather than duplicated; and (2) auditability—every node and edge in $M_{\text{UMM}}$ retains provenance links to its source model(s), allowing downstream validators and human reviewers to trace “where this concept came from.”
Figure 2. Two ways to build the Unified Meta-Model (UMM). Path S1 handles structured sources: when a machine-readable meta-model (for example AUTOSAR, AADL, OPC UA) already exists, ATLAS ingests and normalizes it. Path S2 handles unstructured sources: when only natural-language requirements are available, ATLAS uses an LLM to induce domain entities and relations directly from text. Deployments typically follow one of these two paths, depending on which input is available.
Two-path process for constructing the Unified Meta-Model (UMM). Path S1 applies when a machine-readable meta-model such as AUTOSAR XSD/XMI, AADL, or OPC UA Ecore already exists: ATLAS ingests and normalizes that model using model-to-model transformations. Path S2 applies when no authoritative schema exists and requirements only appear in natural-language documents: ATLAS uses an LLM-based extractor, guided by domain context, to induce entities and relations. Typical deployments choose one of these paths depending on which source type is available.
@@
ATLAS supports two alternative construction paths for the UMM (Figure 2). In practice, deployments usually select one of these paths: either the domain already ships a machine-readable meta-model (Path S1), or it only has natural-language specifications with no canonical XSD/UML (Path S2). Both paths produce the same kind of output: a typed graph $M_{\text{UMM}}$ with provenance annotations. We describe the two paths below.
@@
When an explicit domain meta-model exists, ATLAS reuses an established dual-stage transformation process (DualStageFramework). This process proceeds as follows:
@@
Path S1 therefore produces a UMM instance in which: (i) entities, attributes, and relations reflect the authoritative schema; and (ii) selected behavioral expectations are explicitly attached to those entities. Within ATLAS, this serves as the canonical semantic backbone whenever the domain already provides a formalized meta-model.
@@
Stage 1: Context-Enriched Document Segmentation. A corpus $D=\{d_{1},\ldots,d_{n}\}$ of natural-language specifications is segmented by a learned boundary detector $B$, which identifies sections likely to define entities, attributes, or relationships. For each segment $s_{i}$, ATLAS injects domain context into the prompt (e.g., known entity names, typical field names, role labels from prior segments or from a seed ontology). This grounding step constrains the extractor’s vocabulary and reduces hallucinated concepts.
@@
where $\alpha(c)=\bot$ indicates that $c$ does not match any known node. If $\alpha(c)\neq\bot$, the candidate is treated as an alias or refinement of the matched node, and its provenance metadata is attached to that node. If $\alpha(c)=\bot$, ATLAS tentatively adds a new node or relation into $M_{\text{UMM}}$, tagged with provenance and marked for optional human review when the stakes are high (e.g., safety-critical components). Crucially, at this stage we are only inducing and reconciling domain entities, attributes, and relations; the extraction of constraints are explicitly deferred to the Integrated Constraint Model (ICM) construction in Section 3.3.
@@
Although UMM construction is primarily an integration activity rather than a new theoretical contribution, its runtime characteristics matter for practical deployment. By utilizing union-find with path compression to merge candidate-equivalent nodes, the reconciliation procedure runs in near-linear time relative to the total number of entities across all sources (effectively $\mathcal{O}(\sum|V_{i}|)$). This efficiency allows ATLAS to ingest industrial-scale meta-models with thousands of entities without incurring the performance bottlenecks typical of manual pairwise weaving.
@@
Classical model weaving frameworks (e.g., AMW) rely on manually authored correspondence models, while graph transformation systems (e.g., AGG) require hand-written rules for every pair of source models. Purely LLM-driven extraction has also been explored (chenAutomatedDomainModeling2023), but without grounding the induced concepts in a persistent, provenance-tracked ontology, such approaches risk hallucinated entities and silently inconsistent references. ATLAS differs in three respects:
  • Source Adaptability. Path S1 ingests explicit XSD/XMI-style meta-models through deterministic model transformations (DualStageFramework). Path S2 induces a meta-model from prose using an LLM guided by domain context. Deployers choose whichever path matches their domain reality.
  • Transformation Nature (Fusion vs. Translation). Classical MDE transformations typically strive for bijectivity (or isomorphism) to ensure round-trip engineering capabilities (DualStageFramework; Stevens2010), requiring strict one-to-one mappings between source and target. However, in Path S2, the input (natural language) is inherently redundant, fragmented, and ambiguous, while the target (UMM) must be concise and unified. Consequently, adherence to bijectivity is theoretically ill-suited. ATLAS treats this process not as a translation ($A\leftrightarrow B$) but as a fusion via surjective homomorphisms (formalized by colimits). This shift allows multiple partial textual descriptions to “collapse” into a single semantic entity, rigorously resolving the redundancy and fuzzy alignments that traditional bijective transformations cannot handle.
@@
  • Scalability. Because reconciliation proceeds via incremental union rather than exhaustive pairwise weaving, ATLAS scales to realistic industrial domains, as detailed in the Scalability Note (Section 3.2).
In summary, UMM construction is the step that turns messy, heterogeneous domain knowledge into a single, provenance-aware semantic backbone. All later phases of ATLAS—constraint modeling, generation-time enforcement, post-generation validation, and audit-guided repair—assume this backbone as the shared point of truth.
@@ 3.3. Constraint Extraction and Integrated Constraint Model (ICM)
The Unified Meta-Model (UMM) provides ATLAS with a single, auditable semantic backbone. The Integrated Constraint Model (ICM) builds on that backbone by collecting, normalizing, and compiling the domain rules that generated artifacts must satisfy. These rules range from low-level structural requirements (e.g., mandatory fields, cardinalities) to high-level semantic and logical invariants (e.g., “an OperationInvokedEvent must reference an existing operation”), and they originate from both machine-readable meta-models and natural-language specifications. This subsection formalizes how such constraints are represented, how they are extracted through two complementary channels, how they are grounded in the UMM, and how they are compiled into executable Layer-1 (L1) and Layer-2 (L2) validators for generation-time enforcement and post-generation verification.
@@
Different inputs to ATLAS (e.g., AUTOSAR XSD/XMI schemas vs. regulatory PDFs) induce different constraint spaces. Before integration, these must be semantically aligned with the UMM so that downstream validators can interpret them consistently.
@@
L2 Validator Instantiation and Multi-Modal Enforcement.
The abstract semantic and logic validators $\mathcal{V}_{\mathrm{sem}}$ and $\mathcal{V}_{\mathrm{log}}$ in $\mathcal{I}$ support a multi-modal instantiation strategy to balance theoretical rigor with engineering scalability. In this work, we focus on the formal modality, instantiating semantic validators using SHACL (Shapes Constraint Language) (knublauchSHACLShapesConstraint2017) and logic validators using SMT (Satisfiability Modulo Theories) solvers (demouraZ3EfficientSMT2008). SHACL naturally captures cross-reference semantics in XML/RDF-based artifacts, while SMT efficiently handles bounded numeric reasoning and temporal constraints. This choice prioritizes machine-checkable auditability (“Proof-Carrying”) suitable for safety certification.
However, the architecture also supports a programmatic modality for industrial-scale deployment. By leveraging the strongly-typed model APIs generated via Path S1 (Section 3.2), $\mathcal{C}_{\mathrm{sem}}$ can be compiled into executable validation scripts (e.g., Java/Python assertions). To ensure these scripts are syntactically and semantically valid, we employ type-aware constrained decoding: the LLM’s generation of validation code is strictly masked by the symbol tables of the S2D2-generated APIs. This prevents the model from hallucinating non-existent method calls, effectively creating a “compiler-in-the-loop” generation process. While formal validators provide mathematical completeness for critical safety properties, these verifiable programmatic validators offer superior performance for checking thousands of routine business rules. Regardless of the modality, the key architectural requirement is that validators produce machine-readable certificates $\pi_{\text{sem}}$ and $\pi_{\text{logic}}$ with precise violation localization to drive Audit-Guided Repair.
@@
Channel 2: Inductive Extraction from Natural-Language Specifications. Many industrial domains encode mission-critical rules in prose rather than structured schemas: AUTOSAR PDFs, regulatory guidance, safety standards, legal jurisdiction rules. Channel 2 uses an LLM-based extractor to convert free-form text into structured constraint candidates, following prior work on LLMs as structured information extractors (Dagdelen2024; Lu2022; Wei2020; Jiao2024). Formally,
@@
Figure 3. LLM-based extraction from natural-language specifications. The extractor proposes structured constraint candidates and aligns each candidate to entities in the UMM. Only candidates that align cleanly and pass semantic checks are committed to the Integrated Constraint Model (ICM); the rest are flagged for human review instead of being turned directly into validators.
Illustration of Channel 2 extraction with semantic validation. Specification text is processed by an LLM-based extractor, which proposes structured JSON-like constraint candidates and links each candidate to entities in the Unified Meta-Model (UMM). Only candidates that can be anchored via alpha and that pass semantic compatibility checks are inserted into the Integrated Constraint Model (ICM). Candidates that fail these checks are not compiled directly into L2 validators but are instead flagged for review.
@@
Figure 4. ICM as a multi-sorted algebra layered over the UMM. White nodes denote UMM entities; gray nodes encode semantic and logical constraints extracted from Channel 1 (deductive) and Channel 2 (LLM-assisted). Morphisms relate abstract constraints to class-/attribute-level constraints, enabling cross-layer reasoning and repair prioritization.
Diagram of the Integrated Constraint Model (ICM) represented as a multi-sorted algebra on top of the Unified Meta-Model (UMM). White nodes correspond to UMM entities. Gray nodes capture semantic and logical constraints obtained from Channel 1 (deductive extraction) and Channel 2 (LLM-assisted extraction). Morphisms connect abstract constraints to class- and attribute-level constraints, supporting cross-layer reasoning and guiding which repairs should be prioritized.
@@
Algorithm 1 summarizes how ATLAS assembles the ICM from both structured schemas and unstructured documents. This process is persistent: it populates the long-lived constraint repository (the ICM $\mathcal{I}$), rather than just synthesizing a one-off validator.
@@
In addition to the static ICM, ATLAS supports request-specific constraints that apply only to a single generation run. For example, a user generating one AUTOSAR subsystem may require “all port connections in this subsystem must use compatible data types” without permanently altering the global repository. To accommodate such ad hoc requirements, we synthesize ephemeral L2 validators at runtime:
(11) $\mathrm{Synthesize}^{\mathrm{NLP}}:\mathcal{R}_{\text{dynamic}}\to\mathcal{V}_{\mathrm{L2}}.$
These validators (instantiated as SHACL or SMT in our implementation) are fed directly into L2 verification during CVG (Section 3.4), but they are not persisted in the ICM unless they pass solver-backed validation and are deemed reusable.